OpenStudy (anonymous):

cos2θ = cos^2θ − 1/2

OpenStudy (anonymous):

is this proving \[\cos2\theta=\cos^{2}\theta - \frac{1}{2}\] \[\cos2\theta = \cos(\theta+\theta)=\cos \theta \cos \theta - \sin \theta \sin \theta = \cos^{2}\theta-\sin^{2}\theta\] \[\cos2\theta = \cos^{2}\theta-(1-\cos^{2}\theta)=2\cos^{2}\theta-1\] dividing right-hand side by 2, proves that \[\cos2\theta = \cos^{2}\theta-\frac{1}{2}\]

OpenStudy (anonymous):

Trigonometric Identities used are as follows: \[\cos(\alpha+\beta)=\cos \alpha \cos \beta - \sin \alpha \sin \beta\] \[\sin^{2}\alpha+\cos^{2}\alpha=1\] and \[\cos2\alpha = 2 \cos^{2}\alpha -1\]